Mathematics 131 – Computer Lab Groups

1) Alexandria Andrew and Stephanie Apollon

2) Luke Barr and Oluwaseun Bayode

3) Ha Bui and Amelia Cianci

4) Charles Crawley and Vanessa Dagley

5) Isa Dattilo and Michelle DiBlasi

6) Ashley Domingue and Robert Dudley

7) Kenneth Flanders and Jordan Forester

8) Erika Furtado and Megan Hyland

9) Robert Kerbstat and Katrina Muth

10) Lauren Nakhoul and Kristen Oats

11) Molly O’Keefe and Vondre Ossorio

12) Loyola Pasiewicz and Tara Richards

13) Aaron Rowden and Dominick Sciame

14) Christine Sturdevant and Kristine Sullivan

15) Jason Tartaglione and Michael Thater

16) Elizabeth Umphrey, Victor Vaval, and Gonzalo Vera